We have a Cavendish balance set up with two spheres (one small and one large) that are five centimeters apart of masses m = 1 g and m0 = 500 m, respectively. Suppose these design is placed at a point in space far removed from all other bodies. ,→ What is the acceleration a of the body m, and the acceleration a 0 of the body m0 ? [15] ,→ Is this a constant acceleration problem? Explain why or why not!


1
Expert's answer
2021-05-07T06:41:14-0400
"a=(6.67\cdot10^{-11})\frac{500(0.001)}{0.05^2}=1.3\cdot10^{-8}\frac{m}{s^2}\\
a'=(6.67\cdot10^{-11})\frac{(0.001)}{0.05^2}=2.7\cdot10^{-11}\frac{m}{s^2}"

No. The acceleration will change with distance.